Solve the following system of equations algebraically
4x+3y=12
-2x+y=-16

Found 2 solutions by haterz100, Alan3354:
Answer by haterz100(2)
Answer by Alan3354(69443)

Solve the following system of equations algebraically
4x+3y=12
-2x+y=-16
-------------
Multiply the 2nd eqn by 2, then add them.